# Benchmarking LLM-based Machine Translation on Cultural Awareness

### v2

@@ Appendix A Data Examples
In Table 8, we present a data example from the English-Chinese corpus. Each data point consists of a pair of sentences. We meticulously annotate all cultural-specific items (CSI) within the sentences. For each cultural-specific item, we provide information including its category, country of origin, translations in the target language, descriptions in both the source and target languages, and an explanation. To illustrate the challenges that cultural-specific items pose for current Machine Translation (MT) systems, we provide translations from both Google Translate and ChatGPT for this example. It is noted that both Google and ChatGPT erroneously rendered the Chinese translation of ”Wiener Schnitzel” as ”pork chops” instead of the correct translation, which is ”steak.” This misinterpretation not only misleads Chinese readers but also introduces confusion to the entire sentence.
@@
Target (ZH) 上海炸猪排的做法改良自奥地利国菜
维也纳炸牛排 (Wiener fried steak),而炸猪排
与牛排不同,它显得更加市井。
@@
Translation (ZH) 维也纳炸牛排 (Wiener fried steak)
@@
NLLB 上海风格的炸猪肉切片是从奥地利
国家菜的维也纳施尼切尔(transliteration)改制而成,
LLaMA2 上海炒猪排是一种来自奥地利
牛肉炒肉块 (Beef stir-fried cubes)的改良型,
而炒猪排更像是一道街头小吃而非牛肉炒肉块.
Google Translate 海派炸猪排是奥地利国菜
维也纳炸猪排 (Wiener fried pork chops)
的改良版,炸猪排与其说是牛排,
不如说是街头小吃。
ChatGPT 上海式炸猪排是从奥地利的国菜
维也纳炸猪排 (Wiener fried pork chops)
改编而来,而炸猪排比牛排,
更像是街边食物。
Table 8: A Data Example of the English-Chinese Corpus: In parentheses, we explain what the Chinese translation means.
@@
The mapping table between CSI definition (5 categories in all) and Wikiproject categories [^2] https://en.wikipedia.org/wiki/Wikipedia:WikiProject_Categories (18 categories) are shown in Table 9. Additionally, we provide illustrative examples for each category to clarify the respective meanings. The tool we used for Wikiproject category classification is drafttopic [^3] https://github.com/wikimedia/drafttopic.
@@
Table 9: CSI vs. Wikiproject Mapping Table
@@ Appendix C Wikipedia Parallel Corpus Collection
To collect English-Chinese parallel corpus from Wikipedia, we use the bilingual Wikipedia articles translated through Wikipedia’s content translate tool [^4] https://en.wikipedia.org/wiki/Wikipedia:Content_translation_tool. This tool allows confirmed editors to translate Wikipedia articles from a source language to a target language with a machine translation system. By tracking their editing logs, we obtain the text triples consisting of the original text in a source language, the machine-translated text, and the human post-edited text in a target language. We then use a sentence alignment tool bleu-align [^5] https://github.com/rsennrich/Bleualign Sennrich and Volk (2010) to obtain a sentence-level parallel corpus. Then we use open-source data from OPUS, which includes Wikipediav1.0  [^6] https://opus.nlpl.eu/Wikipedia-v1.0.php for English-French and English-Spanish, as well as Samanantarv0.2  [^7] https://opus.nlpl.eu/Samanantar-v0.2.php for English-Hindi, English-Tamil, and English-Telugu.
@@ Appendix D Data Characteristics
Figure 5: Data characteristics on regions (Outside) and categories (Inside) for the English-Chinese corpus
Culture is intricately linked with specific regions, and its manifestations can exhibit substantial variations across diverse regions and categories. Therefore, our dataset encompasses cultural-specific elements sourced from a wide array of regions and categories, as shown in Figure 5, which shows the distribution of the English-Chinese corpus. This inclusive approach allows us to comprehensively evaluate the performance of machine translation models across a broad spectrum of cultural contexts.
Appendix E Experiment Settings
@@
Appendix F Evaluation Prompts of GPT-4
It’s shown that GPT-4 could be an effective tool for evaluating the generation task’s quality in DPO, we try the same prompt manner with the edition towards understandability of CSIs, the prompt is as follows:
Assuming you’re a Chinese native speaker, which of the following translations does a better job of translating the CSI word “Goubuli” into understandable Chinese words?
¡Source¿
Translation A: ¡A¿
Translation B: ¡B¿
FIRST provide a one-sentence comparison of the two translations, explaining which you prefer and why. SECOND, on a new line, state only ”A” or ”B” to indicate your choice. Your response should use the format:
Comparison: ¡one-sentence comparison and explanation¿
Preferred: ¡”A” or ”B”¿
@@ Appendix G Overall Automatic Evaluation
We first evaluate the translation outputs using traditional automatic metrics such as BLEU Papineni et al. (2002), BLEURT Sellam et al. (2020), and COMET Rei et al. (2020). To be consistent with the evaluation method of NLLB, we calculate spBLEU Goyal et al. (2022) for BLEU scores. In addition to traditional machine translation evaluation metrics, we also use CSI-Match to evaluate the translation quality of CSIs (described in §3). Table 10 shows the comparison results of six MT methods across six language pairs in both directions.
@@
As shown in Table 10, NLLB displays high performance in both directions of translation over six language pairs. Both CSI dictionary incorporation (NLLB-A) and term replacement strategies (NLLB-R) enhance the translation quality of CSI for most language pairs, without significantly compromising the overall sentence translation regarding other metrics. Notably, NLLB-R outperforms other MT systems on CSI-Match, even including LLM-based MT. Interestingly, LLaMA2-7B shows an obvious drop in both traditional evaluation metrics and CSI-Match scores when translating English to three Indian languages and vice versa. One possible explanation is because of the insufficient Indian data during the pre-training of LLaMA2. Both CSI-involving translation strategies are beneficial for LLaMA-based translation. In non-Latin languages (i.e., Chinese, Hindi, Tamil, and Telugu), LLaMA2-A tends to yield better performances, whereas LLaMA2-R performs better in Latin languages (i.e., French and Spanish), which potentially suggests that injecting cultural knowledge through code-switching similar Latin languages works better than distant languages for LLM-based models. Furthermore, we assess the translation performances of ChatGPT and Google Translate. Both MT systems exhibit commendable performance in CSI translation, with Google Translate demonstrating superior translation results. Notably, Google Translate showcases consistent translation abilities, particularly in handling relatively low-resource languages like Tamil and Telugu.
Appendix H Automatic Evaluation of Prompting Strategies on LLaMA2
while on LLaMA2, we employed a 2-shot prompting method, which is different from the previous 8-shot in-context learning (§6.1) due to the LLaMa2-7B input length limit. The understandability of LLaMA-2 is shown in Table 11. 2-shot prompting strategies can hardly improve the understandability of LLaMA2-7B. The complexity of prompts even causes unstable generation results which show worse performance than baseline.
Shot Method CSI-Match Overall-U NT-U
two BI 35.1 16.3 22.2
CT 70.3 16.0 -
CE 35.5 11.1 14.2
SE 35.1 11.8 15.8
SR 37.9 10.2 11.1
Table 11: Evaluation of different strategies by LLaMA2 on English-Chinese translations.
Appendix I Performance Across Regions
Figure 6: Avg.CSI-Match by Regions
@@ Appendix I Performance Across Regions
Culture is often associated with a specific region, and its expressions can vary significantly across different regions and categories. To gain a deeper understanding of the influence of region on CSI translation, we categorized CSI into three groups: CSI originating from countries primarily using the source language, countries predominantly using the target language, and countries utilizing languages other than the source and target languages. On the six groups of English-to-XX translations, we calculated the average CSI-Match values of these three CSI groups respectively, and the results are shown in Figure 6.
@@
Given that target CSIs must have the translation in the target language. Translating target CSIs is akin to back translation. However, when translating the source CSI or other CSIs, the translation may either not exist in the target language or exist with lower word frequency. Consequently, the model is expected to yield better results for the target CSI. Surprisingly, our analysis reveals that most models excel at translating the target CSI back into the target language in Latin languages (i.e., French and Spanish). Notably, Google Translate consistently achieves superior translations across all languages. ChatGPT demonstrate better translation performance in Chinese and Tamil, while LLaMA2 succeeds in Hindi and Tamil for target CSI translation. In contrast, traditional translation models NLLB struggle with all non-Latin languages, failing to outperform the source CSI translation. This suggests that LLMs may possess enhanced learning capabilities for translating culture-related content. However, it is important to note that the current translation performance is not consistently stable.
Appendix J Comparison of Translation Strategies
@@
Figure 7: Comparison of Translation Strategies: The value in each grid represents the winning rate of the method on the x-axis in comparison to the method on the y-axis.
@@
In order to further compare the impact of different translation strategies on understandablity, we analyzed the comparison between different translation strategies based on the ranking results of human evaluation. Specifically, we also rank the different translation strategies used by the MT systems according to the rank of the MT system’s understandablity given by humans, which is shown in Figure 7. It’s shon that the winning rate of descriptions for all methods surpasses 0.5, and the winning rate of substitutions, excluding descriptions, significantly exceeds 0.5. This implies that translations employing these two strategies are generally deemed more comprehensible by human annotators. Moreover, Literal Translation outperforms Transliteration, highlighting that transliteration may diminish the clarity of CSI in translation compared to a literal approach. Notably, the winning rate of copying for both Literal Translation and Transliteration hovers around 50%, indicating that these two methods may introduce confusion, and their readability underperforms directly copying the original word.

### v3

@@ Appendix A Data Examples
In Table 8, we present a data example from the English-Chinese corpus. Each data point consists of a pair of sentences. We meticulously annotate all culture-specific items (CSI) within the sentences. For each culture-specific item, we provide information including its category, country of origin, translations in the target language, descriptions in both the source and target languages, and an explanation. To illustrate the challenges that cultural-specific items pose for current Machine Translation (MT) systems, we provide translations from both Google Translate and ChatGPT for this example. It is noted that both Google and ChatGPT erroneously rendered the Chinese translation of "Wiener Schnitzel" as "pork chops" instead of the correct translation, which is "steak". This misinterpretation not only misleads Chinese readers but also introduces confusion to the entire sentence, whose meaning is “The Shanghai-style pork chops are a twist on Austria’s national dish, Wiener fried pork chops, which are more street food than steak.
@@
Target (ZH) {CJK}UTF8gbsn上海炸猪排的做法改良自奥地利国菜
{CJK}UTF8gbsn维也纳炸牛排 (Wiener fried steak),而炸猪排
{CJK}UTF8gbsn与牛排不同,它显得更加市井。
@@
Translation (ZH) {CJK}UTF8gbsn维也纳炸牛排 (Wiener fried steak)
@@
NLLB {CJK}UTF8gbsn上海风格的炸猪肉切片是从奥地利
{CJK}UTF8gbsn国家菜的维也纳施尼切尔(transliteration)改制而成,
LLaMA2 {CJK}UTF8gbsn上海炒猪排是一种来自奥地利
{CJK}UTF8gbsn的牛肉炒肉块 (Beef stir-fried cubes)的改良型,
{CJK}UTF8gbsn而炒猪排更像是一道街头小吃而非牛肉炒肉块.
Google Translate {CJK}UTF8gbsn海派炸猪排是奥地利国菜
{CJK}UTF8gbsn 维也纳炸猪排 (Wiener fried pork chops)
{CJK}UTF8gbsn的改良版,炸猪排与其说是牛排,
{CJK}UTF8gbsn不如说是街头小吃。
ChatGPT {CJK}UTF8gbsn上海式炸猪排是从奥地利的国菜
{CJK}UTF8gbsn 维也纳炸猪排 (Wiener fried pork chops)
{CJK}UTF8gbsn改编而来,而炸猪排比牛排,
{CJK}UTF8gbsn更像是街边食物。
Table 8: A Data Example in the English-Chinese Corpus: in parentheses, we explain what the Chinese translation means.
@@
Table 9: CSI vs. Wikiproject mapping table.
The mapping table between CSI definitions (5 categories in total) and Wikiproject categories [^2] https://en.wikipedia.org/wiki/Wikipedia:WikiProject_Categories (18 categories) are shown in Table 9. Additionally, we provide examples for each category to clarify the respective meanings. The tool we used for Wikiproject category classification is drafttopic [^3] https://github.com/wikimedia/drafttopic.
@@ Appendix C Wikipedia Parallel Corpus Collection
To collect the English-Chinese parallel corpus from Wikipedia, we use the bilingual Wikipedia articles translated through Wikipedia’s content translate tool [^4] https://en.wikipedia.org/wiki/Wikipedia:Content_translation_tool. This tool allows confirmed editors to translate Wikipedia articles from the source language to a target language with a machine translation system. By tracking their editing logs, we obtain the text triples consisting of the original text in a source language, the machine-translated text, and the human post-edited text in the target language. We then use a sentence alignment tool bleu-align [^5] https://github.com/rsennrich/Bleualign Sennrich and Volk (2010) to obtain a sentence-level parallel corpus. To obtain more language pairs, we reuse open-source data from OPUS, which includes Wikipediav1.0  [^6] https://opus.nlpl.eu/Wikipedia-v1.0.php for English-French and English-Spanish, as well as Samanantarv0.2  [^7] https://opus.nlpl.eu/Samanantar-v0.2.php for English-Hindi, English-Tamil, and English-Telugu.
@@ Appendix D Data Characteristics
Data Statistics
Table 2 shows the statistics of our parallel corpora for the evaluation of MT systems on six language pairs. Particularly, for each language pair, we count the total number of detected CSIs by CSIs Counts and the number of unique CSIs by CSIs Types. It’s noted that not all the CSIs have translations on Wikidata, so we determine the number of CSIs containing translations in WikiData by CSI Translations. Considering that many CSIs only exist within a specific culture group, which can’t be located in the parallel corpus, CSIs that don’t have a translation in other languages should take a higher proportion in the real-world corpus than in our dataset.
Data Diversity
Culture is intricately linked to specific regions, and its manifestations can exhibit substantial variations across diverse regions and categories. Therefore, our dataset encompasses culturally specific items sourced from a wide array of regions and categories. Figure 5 shows the distribution of categories. Specifically, we mapped 18 Wikiproject categories into 5 culture categories. Since there is no Wikiproject category matching the CSI category gestures and habits, we excluded this label from further consideration. Regarding the regions, we show the top 15 origin countries in our dataset in Figure 6. Among these regions, CSIs originating from English-speaking countries (e.g., the United States and the United Kingdom) have the highest representation. This is because we conduct entity-linking on the English source texts, resulting in a predominance of CSIs from English-speaking countries. However, the entity linking tool SLING [^8] https://github.com/google/sling is multilingual, making it feasible to use our pipeline to include more CSIs from non-English speaking countries. This inclusive approach allows us to comprehensively evaluate the performance of machine translation models across a broad spectrum of cultural contexts.
Figure 5: Category distribution on categories.
Figure 6: Data characteristics on regions.
Appendix E Evaluation Prompts of GPT-4o
It has been shown that GPT-4 can be an effective tool for evaluating the quality of generation tasks in DPO Rafailov et al. (2023). We apply a similar prompt method for the pragmatic translation quality evaluation of CSIs. The prompt is as follows, with the system output and reference randomly shuffled into choices A and B:
Assuming you’re a Chinese native speaker, which of the following translations has a more understandable translation in Chinese of following culture-specific item: “Goubuli”? Please only compare the itemś translation by ignoring the translation quality and length of the whole sentence.
<Source>
Translation A: <A>
Translation B: <B>
FIRST, provide a one-sentence comparison of the two translation, explaining which you prefer and why. SECOND, on a new line, if the translations of cultural-specific items: “Goubuli” in "A" and "B" are different, state "A" or "B" to indicate your choice, otherwise, use "C" to indicate your choice. Your response should use the format:
Comparison: <one-sentence comparison and explanation>
Preferred: <"A" or "B" or "C">
For the human evaluation, we also use the same prompt as instructions to align the human evaluations with GPT-4o evaluations.
Appendix F Experiment Settings
@@ Appendix G Overall Automatic Evaluation
We evaluate the translation outputs using traditional automatic metrics such as BLEU Papineni et al. (2002b), BLEURT Sellam et al. (2020), and COMET Rei et al. (2020b). To be consistent with the evaluation method of NLLB, we calculate spBLEU Goyal et al. (2022) for BLEU scores. In addition to traditional machine translation evaluation metrics, we also use CSI-Match to evaluate the translation quality of CSIs (described in §4). Table 10 shows the results of eight MT systems across six language pairs in two directions.
@@
As shown in Table 10, both CSI dictionary incorporation (NLLB-A) and term replacement strategies (NLLB-R) enhance the translation quality of CSI for most language pairs, without significantly compromising the overall sentence translation regarding other metrics. Notably, NLLB-R outperforms other MT systems on CSI-Match, even including LLM-based MT. Interestingly, LLaMA2-7B shows an obvious drop in both traditional evaluation metrics and CSI-Match scores when translating English to three Indian languages and vice versa. One possible explanation is because of the insufficient Indian data during the pre-training of LLaMA2. Both CSI-involving translation strategies are beneficial for LLaMA-based translation. In non-Romance languages (i.e., Chinese, Hindi, Tamil, and Telugu), LLaMA2-A tends to yield better performances, whereas LLaMA2-R performs better in Romance languages (i.e., French and Spanish), which potentially suggests that injecting cultural knowledge through code-switching similar Romance languages works better than distant languages for LLM-based models. Furthermore, we assess the translation performances of ChatGPT and Google Translate. Both MT systems exhibit commendable performance in CSI translation, with Google Translate demonstrating superior translation results. Notably, Google Translate showcases consistent translation abilities, particularly in handling relatively low-resource languages like Tamil and Telugu.
Appendix H PTA Evaluation Results Across Languages
We evaluate the PTA of two more language pairs. The evaluation result is shown in Figure 7. As with CSI-match on these two languages, the PTA performances of the 4 MT systems are pretty close. However, GPT-3.5 still shows superior performance on PTA compared to NMTs, indicating that GPT-3.5 has better capabilities to generate free translations for CSIs which can be easily understood by native speakers in the target culture.
Figure 7: PTA results on En-Fr and En-Es translations.
Appendix I Generation Data Examples of Non-translation CSIs
Table 11 shows the results of the four prompting strategies on the CSI “milk toast” from English to Chinese, which has no known translations. Under the BI strategy, GPT-3.5 translates the American breakfast dish as “toast”, failing to capture the defining feature of the dish, which is that it is soaked in milk. CT similarly fails, yielding the term “milk bread”. “Milk” as an adjective does not adequately describe the dish, and "bread" no longer specifies the toasted aspect. The former issue likewise arises with CE, a literal translation of “milk toast”. In contrast, using SE, GPT-3.5 integrates the CSI explanation into the translation, freely translating the term as “toasted bread soaked in milk”. This makes it easier for Chinese readers to understand the meaning of "milk toast", as is reflected in the ratings of GPT-4 and the human annotator.
Strategy Outputs PTA of GPT-4o PTA of Human
BI {CJK}UTF8gbsn该书中最终包含了1850种食谱,其中有烤面包(toast) Lose Lose
CT {CJK}UTF8gbsn该书最终收录了1850个食谱,比如牛奶面包(milk bread) Lose Lose
CE {CJK}UTF8gbsn这本书最终包含了1850个食谱,其中有牛奶土司(literal translation)。 Same Same
SE {CJK}UTF8gbsn该书最终包含了1850个配方,其中有烤面包浸牛奶(toasted bread soaked in milk)。 Win Win
Source The book eventually contained 1,850 recipes including milk toast.
Reference {CJK}UTF8gbsn书中收录了1850个食谱,其中有牛奶土司 (literal Translation)
Knowledge Translation: No existing Chinese translations
Explanation: the entity is a breakfast dish consisting of toasted bread in warm milk, typically with sugar and butter…
Table 11: Non-translation CSI output examples of prompting strategy on GPT-3.5 and a source-reference sentence pair with cultural knowledge for En-Zh translations.
Appendix J Generation Examples of LLaMA
Strategy Outputs PTA of GPT-4o PTA of Human
BI {CJK}UTF8gbsn爱尔兰早餐卷的制作方式类似于一顿早餐休闲袋(breakfast relaxing bag) Lose Lose
CT {CJK}UTF8gbsn爱尔兰早饭卷饼是与墨西哥卷饼(copy the dictionary)准确相似的。 Win Win
CE {CJK}UTF8gbsn爱尔兰早餐卷的制作方式与美国西南部的早餐卷(American southwest breakfast roll)一样。 Win Win
SE {CJK}UTF8gbsn爱尔兰的早餐卷是和早餐壳(breakfast shell)。 Win Win
Source The breakfast roll of Ireland is prepared similarly to a breakfast burrito.
Reference {CJK}UTF8gbsn爱尔兰的早餐面包卷制作方法亦类似于此 (Replace the term with it)
Knowledge Translation:{CJK}UTF8gbsn墨西哥卷饼
Explanation: The entity is a dish in Mexican and Tex-Mex cuisine, consisting of
a flour tortilla wrapped into a sealed cylindrical shape around various ingredients…
Table 12: CSI output examples of prompting strategy on LLaMA and a source-reference sentence pair with cultural knowledge for En-Zh translations.
Table 12 shows the results of four prompting strategies on the CSI “burrito” from English to Chinese, defined as a "flour tortilla wrapped into a sealed cylindrical shape around various ingredients." Under the BI and CE strategies, LLaMA translates it as “bag” and “shell” respectively, failing to capture the essential feature of the dish, which is its rolled shape. The CT strategy successfully copies the dictionary translation. Interestingly, CE freely translates the word into "American southwest breakfast roll," accurately describing the food’s shape. Additionally, CE prompts LLaMA to leverage related cultural knowledge to include the region description in the translation of the CSI.
Appendix K Performance Across Regions
Figure 8: Avg.CSI-Match by regions.
@@ Appendix I Generation Data Examples of Non-translation CSIs
Culture is often associated with a specific region, and its expressions can vary significantly across different regions and categories. To gain a deeper understanding of the influence of region on CSI translation, we categorized CSI into three groups: CSI originating from countries primarily using the source language, countries predominantly using the target language, and countries utilizing languages other than the source and target languages. In the six groups of English-to-XX translations, we calculated the average CSI-Match values of these three CSI groups respectively, shown in Figure 8.
@@
Given that target CSIs must have the translation in the target language, translating target CSIs is akin to back translation. However, when translating the source CSI or other CSIs, the translation may either not exist in the target language or exist with lower word frequency. Consequently, the model is expected to yield better results for the target CSI. Surprisingly, our analysis reveals that most models excel at translating the target CSI back into the target language in Romance languages (i.e., French and Spanish). Notably, Google Translate consistently achieves superior translations across all languages. ChatGPT demonstrates better translation performance in Chinese and Tamil, while LLaMA2 succeeds in Hindi and Tamil for target CSI translation. In contrast, traditional translation models NLLB struggle with all non-Romance languages, failing to outperform the source CSI translation. This suggests that LLMs may possess enhanced learning capabilities for translating culture-related content. However, it is important to note that the current translation performance is not consistently stable.
Appendix L Comparison of Translation Strategies
Figure 9: Percentage of translation strategies.
Translation Strategies
To explore potential factors benefiting pragmatic translation quality, we let a human annotator examine the models’ translation strategies. We categorize the translation strategies of CSIs based on prior translation theories Newmark (1988); Persson (2015). These theories define different categorizations of strategies to improve the comprehensibility of CSIs while maintaining cultural integrity. We select 4 strategies that are common to our dataset. They’re 1) Transliteration that phonetically translates source CSIs; 2) Literal translation that directly translates word-by-word; 3) Description that integrates CSI descriptions of the CSIs into the translation; 4) Substitution that replace source CSIs by a semantically equivalent item in the target language; 5) Copy that directly copies the source language of CSI into the target language; 6) Wrong that indicates entirely incorrect translations; and 7) Other that employs other strategies in translation. Figure 9 shows the ratio of each strategy in four MT systems. We find that models with higher PTA (e.g., ChatGPT and Google translate) use description and substitution at a significantly higher rate, indicating that these two strategies help improve the understanding of CSI for target-language speakers. Notably, LLaMA2 incorporates a higher frequency of substitution and description methods compared to traditional NLLB. However, this increased diversity in translation output comes at the cost of reduced stability in the outputs. As a result, LLaMA2 tends to yield more inaccurate translations, whereas NLLB relies more on Literal Translation and Transliteration to translate CSIs.
@@
Figure 10: Comparison of Translation Strategies: The value in each grid represents the win rate of the method on the x-axis in comparison to the method on the y-axis.
@@
In order to further compare the impact of different translation strategies on pragmatic translation quality, we analyzed the comparison between different translation strategies based on the ranking results of human evaluation. Specifically, we also rank the different translation strategies used by the MT systems according to the rank of the MT system’s comprehensibility given by humans, which is shown in Figure 10. It’s shown that the win rate of descriptions for all methods surpasses 0.5, and the win rate of substitutions, excluding descriptions, significantly exceeds 0.5. This implies that translations employing these two strategies are generally deemed more comprehensible by human annotators. Moreover, Literal Translation outperforms Transliteration, highlighting that transliteration may diminish the clarity of CSI in translation compared to a literal approach. Notably, the win rate of copying for both Literal Translation and Transliteration hovers around 50%, indicating that these two methods may introduce confusion, and their readability underperforms directly copying the original word.